OpenStudy (anonymous):

Solve the equation 2a = 3a + 4c for a

OpenStudy (shane_b):

2a=3a+4c 2a-3a=4c -a=4c a=-4c

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[0 = 3a - 2a + 4c\] \[0 = a + 4c\] \[a = -4c\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Don't thank us twice.. say thanks to Shane B..

OpenStudy (anonymous):